Grade 11General Physics

An X-ray tube produces a continuous spectrum of radiation with its short wavelength end at 0.45 Å. What is the maximum energy of a photon in the radiation?

Jayant Kumar
12 Years agoGrade 11

1 Answer

Vikas TU
6 Years ago
Dear student 
λ=0.45×10^−10m
maximum energy is E=hc/λ=27.6×10^3
eV=27.6keV
…................................